You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121205 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 4-5

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwars
Instrument: LUCI1

Summary

Tests of the rotator during the evening. Dave, Jhon and Geno supervise and test the telescope performances.
Bouyancy problems are solved before to open the dome.
At the twilight clear sky with thin cirrus only at the orizon.
The seeing at the beginning of the night is 0.9".
During the Cresci_28A target integration, both the script and the WFs paused. The actual problem originates from the Guider hardware. Data are NOT affected by this issue.
During the nigh thin cirrus passing. Seeing 0.6"0.7".
Low wind during the entire night.
At the end of the night we have thick clouds passing

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_28A DSF2237a K, 210_zJHK 40m Tell. Integration interrupt. Guider fault
Pentericci_02 1UDS HKspec, 200_H+K 2h20m Tell+flat+arc
Andreon_01 JKCS041 HKspec, 200_H+K 50m Tell+flat+arc
LaFranca_04 Mrk417 zJspec, 200_H+K 45m Tell+flat+arc. COMPLETED
LaFranca_10 NGC4395 zJspec, 200_H+K 45m Tell+flat+arc COMPLETED

11.50 The T.O. open the Dome

13.00 We slew to the telluric Cresci_28A
  • executeLuciscript.sh tell_campoA_acq.txt luci20121205.0015-0017

13.06 executeLuciscript.sh tell_int_campoA_K.txt luci20121205.0018-0025

13.10 executeLuciscript.sh acq_campoA.txt luci20121205.0026-0030

13.33 executeLuciscript.sh sci_campoA_K_1.txt luci20121205.0031-0038

02.26 Problem The script has paused. The WFs paused. The T.O. calls John Hill.
  • We have to reacquire
  • The S.A. ensures that the problems do not affect the data already taken

03.21 executeLuciscript.sh acq_campoA.txt luci20121205.0039-0044
  • The seeing is 0.6"

03.48 Time lost to reacquire Cresci_28A, stars in the reference boxes are too faint possibly because the target is reaching the airmass limit.
  • We slew to Pentericci_02.
  • executeLuciscript.sh pente1_telluric.acq luci20121205.0045-0047

04.00 executeLuciscript.sh pente1_telluric_mirror.spec luci20121205.0048-0050

04.06 executeLuciscript.sh pente1.acq luci20121205.0051-0054

04.24 executeLuciscript.sh pente1_HK_1.spec luci20121205.0055-0066

05.33 executeLuciscript.sh pente1_HK_2.spec luci20121205.0067-0078

06.38 executeLuciscript.sh pente1_HK_3.spec luci20121205.0079-0082

07:02 Slew to Andreon
  • executeLuciscript.sh andreon_mos.acq luci20121205.0083-0085

07:18 executeLuciscript.sh andreon_HK_1.spec luci20121205.0086-0095

08.13 We slew to the telluric for Andreon_01
  • executeLuciscript.sh andreon_telluric_mir.acq luci20121205.0096-0098

08.24 executeLuciscript.sh andreon_telluric_mir.spec luci20121205.0099-0106
  • The spectrum is in the first slit, bur the offsets are wrong. We decide to reacquire the spectrum by hand in a blue and in a red slit.

08.32 executeLuciscript.sh andreon_telluric_mir.acq luci20121205.0107-109

08.41 executeLuciscript.sh andreon_telluric_mir.spec luci20121205.0110
  • standard star in the red slit

08.44 executeLuciscript.sh andreon_telluric_mir.acq luci20121205.0111

08.50 executeLuciscript.sh andreon_telluric_mir.spec luci20121205.0112
  • standard star in the blue slit

08.54 We slew to LaFranca_04
  • executeLuciscript.sh STDA4_1_J.acq
  • Wrong camera, we need to go the Zenith.

09.00 executeLuciscript.sh STDA4_1_J.acq
  • LUCI software exits the acquisition script, but the preset is actually running

09.07 executeLuciscript.sh STDA4_1_J.acq luci20121205.0113-0119

09.15 executeLuciscript.sh STDA4_1_200_zJspec.spec luci20121205.0120-0121
  • we were low with the previous target so we slew to a bright star close to Mrk417 to collimate

09.22 executeLuciscript.sh A4_1_J.acq luci20121205.0122-0126
  • some pattern in the spectral frame due to condensation in the camera.
  • collimation problem
  • we were low with the previous target so we slew to a bright star close to Mrk417 to collimate

09:36 executeLuciscript.sh A4_1_200_zJspec.spec luci20121205.0127-0134

10.29 We Slew to target LaFranca_A11/B2
  • executeLuciscript.sh STDA11_1_J.acq luci20121205.0135-0138
  • clouds are passing

10.37 executeLuciscript.sh STDA11_1_200_zJspec.spec luci20121205.0139-0140

10.42 executeLuciscript.sh A11_1_J.acq luci20121205.0141-0143
  • It is not possible to guide due to a combination of passing clouds and proximity of the moon
  • We point a bright star to collimate

11.17 THICK clouds We have to stop for a while with the scientific observations because it is not possible to collimate even with very bright stars.

11:30 After clouds
  • executeLuciscript.sh OB_LaFranca_A10_acq.txt luci20121205.0144-0149

11:43 executeLuciscript.sh OB_LaFranca_A10_spe.txt luci20121205.0150-0157
  • seeing 0.5"

12.37 executeLuciscript.sh STDA10_1_J.acq luci20121205.0158-0162

12.45 executeLuciscript.sh STDA10_1_200_zJspec.spec luci20121205.0163-0164

12.52 The T.O. statrs to close the dome

13.05 We start calibrations
  • executeLuciscript.sh All_1_200_zJspec.flat luci20121205.0165-0204
  • executeLuciscript.sh andreon_HK.flat luci20121205.0205-0208
  • executeLuciscript.sh pente1_HK.flat luci20121205.0209-0212
  • executeLuciscript.sh All_1_210_K.flat luci20121205.0213-0252
  • executeLuciscript.sh All_1_210_K.arc luci20121205.0253-0270
  • executeLuciscript.sh pente1_HK.arc luci20121205.0271-0274
  • executeLuciscript.sh All_1_200_zJspec.arc luci20121205.0279-0290

-- VincenzoTesta - 04 Dec 2012
Topic revision: r8 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback